# Hybrid Inverse Reinforcement Learning

### v1

@@ 2 Related Work
Hybrid Training for Inverse Reinforcement Learning. Perhaps the most similar approaches to our model-free HyPE algorithm are the SQIL approach of Reddy et al. (2019) and the AdRIL approach of Swamy et al. (2021). Both approaches use data from both the learner and the expert during policy updates by sampling from two separate replay buffers. However, neither of these works rigorously addresses the effect of using off-policy data in the policy optimization subroutine of inverse RL. We provide a general reduction that specifies the properties required for doing so while preserving performance guarantees. We are unaware of any model-based hybrid IRL approaches like HyPER.
@@ 3.1 A Game-Theoretic Perspective on Inverse RL
We consider a finite-horizon Markov Decision Process (MDP) (Puterman, 2014) parameterized by $\langle\mathcal{S},\mathcal{A},\mathcal{T},r,H\rangle$ where $\mathcal{S}$, $\mathcal{A}$ are the state and action spaces, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ is the transition operator, $r:\mathcal{S}\times\mathcal{A}\rightarrow[-1,1]$ is the reward function, and $H$ is the horizon. In the inverse RL setup, we see trajectories generated by an expert policy $\pi_{E}:\mathcal{S}\rightarrow\Delta(\mathcal{A})$, but do not know the reward function. Our goal is to find a policy that, no matter what reward function we are evaluated under, performs as well as the expert. We cast this problem as a zero-sum game between a policy player and an adversary that tries to pick out differences between expert and learner policies (Swamy et al., 2021). More formally, we optimize over policies $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})\in\Pi$ and reward functions $f:\mathcal{S}\times\mathcal{A}\rightarrow[-1,1]\in\mathcal{F}_{r}$. We use $\xi=(s_{1},a_{1},r_{1},\ldots,s_{H},a_{H},r_{H})$ to denote the trajectory generated by some policy. For theoretical simplicity, we assume that our strategy spaces ($\Pi$ and $\mathcal{F}_{r}$) are convex and compact, that $\mathcal{F}_{r}$ is closed under negation, and that $r\in\mathcal{F}_{r},\pi_{E}\in\Pi$. Using $J(\pi,\hat{r})=\mathbb{E}_{\xi\sim\pi}[\sum_{h=1}^{H}\hat{r}(s_{h},a_{h})]$ to denote the value of policy $\pi$ under reward function $\hat{r}$, we can express our objective as
@@
Indeed, one may need to explore the entire state space for the optimal policy when given an arbitrary reward function. However, in inverse RL we explicitly choose reward functions that rate the expert higher than the history of learner policies. Thus, the fact that the standard recipe for both both primal and dual IRL completely ignore the expert demonstrations during policy optimization seems suboptimal. This begs the question: how can we give the learner examples of expert behavior during policy optimization to reduce the amount of exploration required to find good states? We now discuss the core principle that underlies multiple ways to do so while preserving performance guarantees.
@@ 4 Experiments
Figure 4: We see HyPER and HyPE achieve the highest reward on the Mujoco locomotion benchmark. Further, the performance gap increases with the difficulty of the environment (i.e. how far right a plot is in the above figure). We run all model-free algorithms for 1mil environment steps. Due to the higher interaction efficiency of model-based approaches, we only run HyPER for 150k environment steps, after which the last reward is extended horizontally across. We compute standard error across 5 seeds for HyPER, and across 10 seeds for all other algorithms.
@@
We implement HyPE by updating the policy and critic networks in Soft Actor Critic (Haarnoja et al., 2018) with expert and learner samples. We implement HyPER by running model-based policy optimization (Janner et al., 2019) and resetting to expert states in the learned model. No reward information is provided in either case, so we also train a discriminator network. We compare against four baselines: BC (behavioral cloning, Pomerleau (1988)), IQLearn (Garg et al., 2021), MM (a baseline moment-matching algorithm that uses an integral probability metric instead of Jensen-Shannon divergence as suggested by Swamy et al. (2022)), BC-Reg (MM with an added Mean-Square Error Loss on the actor update), and FILTER (IRL with resets to expert states,  Swamy et al. (2023)). Appendix B includes additional implementation details and hyperparameters. We release the code we used for all of our experiments at https://github.com/jren03/garage.
Mujoco Experiments. On the Mujoco locomotion benchmark environments (Brockman et al., 2016), all learners are provided with 64 demonstration trajectories from an RL expert to mitigate finite sample concerns. To harden the problem, noise is injected into the environment in the following manner: with probability $p_{tremble}$, a random action is executed instead of the one chosen by the policy. As seen in Figure 4, HyPE and HyPER converge much quicker to expert performance compared to baselines, and the gap increases as environments get more difficult (further right in the figure). While our algorithms have the same worst-case performance bound as FILTER, HyPE and HyPER show to be empirically more sample efficient. We hypothesize this is because HyPE and HyPER use expert states and actions, while FILTER only uses expert states. We find these results to be particularly exciting, as both algorithms show competitive performance and sample efficiency without needing expert resets. Finally, IQLearn fails to match BC performance on some environments, and shows signs of unstable training on others. We suspect this is due to the need to model environment stochasticity implicitly.
@@
5 Impact Statement
@@
6 Acknowledgments
ZSW is supported in part by the NSF FAI Award #1939606, a Google Faculty Research Award, a J.P. Morgan Faculty Award, a Facebook Research Award, an Okawa Foundation Research Grant, and a Mozilla Research Grant. GS would like to thank Yuda Song for many helpful edits and being a hybrid RL oracle throughout the course of this project.
@@ A.1 Miscellaneous Lemmas
$+\mathbb{E}_{\xi\sim\pi_{E},M^{\star}}\left[\sum_{h}^{H}\mathbb{E}_{\begin{subarray}{c}s^{\star}_{h+1}\sim M^{\star}\\ a\sim\pi(s_{h+1}^{\star})\end{subarray}}[\hat{Q}_{r}(s^{\star}_{h+1},a)]-\mathbb{E}_{\begin{subarray}{c}\hat{s}_{h+1}\sim\hat{M}\\ a\sim\pi(\hat{s}_{h+1})\end{subarray}}[\hat{Q}_{r}(\hat{s}_{h+1},a)]\right]$ (16)
$+\mathbb{E}_{\xi\sim\pi,M^{\star}}\left[\sum_{h}^{H}\mathbb{E}_{\begin{subarray}{c}\hat{s}_{h+1}\sim\hat{M}\\ a\sim\pi(\hat{s}_{h+1})\end{subarray}}[\hat{Q}_{r}(\hat{s}_{h+1},a)]-\mathbb{E}_{\begin{subarray}{c}s^{\star}_{h+1}\sim M^{\star}\\ a\sim\pi(s_{h+1}^{\star})\end{subarray}}[\hat{Q}_{r}(s^{\star}_{h+1},a)]\right].$ (21)
@@
$=\mathbb{E}_{\xi\sim\pi_{E},M^{\star}}[\sum_{h=0}^{H}(\mathbb{E}_{\begin{subarray}{c}s^{\star}_{h+1}\sim M^{\star}(s_{h},a_{h})\\ a\sim\pi_{E}(s^{\star}_{h+1})\end{subarray}}[\hat{Q}_{r}(s^{\star}_{h+1},a)]-\mathbb{E}_{\begin{subarray}{c}s^{\star}_{h+1}\sim M^{\star}(s_{h},a_{h})\\ a\sim\pi(s^{\star}_{h+1})\end{subarray}}[\hat{Q}_{r}(s^{\star}_{h+1},a)])$
$-\sum_{h=0}^{H}(\mathbb{E}_{\begin{subarray}{c}\hat{s}_{h+1}\sim\hat{M}(s_{h},a_{h})\\ a\sim\pi_{E}(\hat{s}_{h+1})\end{subarray}}[\hat{Q}_{r}(\hat{s}_{h+1},a)]-\mathbb{E}_{\begin{subarray}{c}s^{\star}_{h+1}\sim M^{\star}(s_{h},a_{h})\\ a\sim\pi(s^{\star}_{h+1})\end{subarray}}[\hat{Q}_{r}(s^{\star}_{h+1},a)])].$
@@
(T1) + (T3) $=\mathbb{E}_{\xi\sim\pi_{E},M^{\star}}[\sum_{h=0}^{H}(\mathbb{E}_{\begin{subarray}{c}a\sim\pi_{E}(s_{h})\end{subarray}}[\hat{Q}_{r}(s^{\star}_{h+1},a)]-\mathbb{E}_{\begin{subarray}{c}a\sim\pi(s_{h})\end{subarray}}[\hat{Q}_{r}(s^{\star}_{h+1},a)])$
$+\sum_{h=0}^{H}(\mathbb{E}_{\begin{subarray}{c}s^{\star}_{h+1}\sim M^{\star}(s_{h},a_{h})\\ a\sim\pi(s^{\star}_{h+1})\end{subarray}}[\hat{Q}_{r}(s^{\star}_{h+1},a)]-\mathbb{E}_{\begin{subarray}{c}\hat{s}_{h+1}\sim\hat{M}(s_{h},a_{h})\\ a\sim\pi_{E}(\hat{s}_{h+1})\end{subarray}}[\hat{Q}_{r}(\hat{s}_{h+1},a)])].$
@@
[…]
@@ B.1 Mujoco Tasks
Discriminator. For our discriminator, we start with a learning rate of $48e-$ 4 and decay it linearly over outer-loop iterations. For all model-free Mujoco experiments, we update the discriminator every 10,000 actor steps. For all model-based Mujoco experiments, we update the discriminator every 2,000 actor steps.
@@
HyPER. For HyPER, we use the implementation from Pineda et al. (2021) with modifications on the actor update according to Vemula et al. (2023), and turn on the entropy bonus. The model is provided the same demonstration dataset of 64,000 transition tuples as model-free experiments. We use an ensemble of discriminators equal to the model ensemble size, and take the minimum value of the ensemble each forward pass. Further, we find that clipping the discriminator output and adding the same weight decay to the model and actor networks helps stabilize performance for some environments. We list the specific hyperparameters used for each environment in Tables 3 to 6.
@@ B.2 D4RL Tasks
HyPE. Both HyPE and HyPE+Resets use the same TD3+BC optimizer and hyperparameters for the actor as MM and FILTER, and is pretrained with 10,000 steps of behavioral cloning. HyPE+Resets uses $alpha=1.0$ to always reset to expert states.

### v2

@@ 2 Related Work
Hybrid Training for Inverse Reinforcement Learning. Perhaps the most similar approaches to our model-free HyPE algorithm are the SQIL approach of Reddy et al. (2019) and the AdRIL approach of Swamy et al. (2021). Both approaches use data from both the learner and the expert during policy updates by sampling from two separate replay buffers. However, neither of these works rigorously addresses the effect of using off-policy data in the policy optimization subroutine of inverse RL. We provide a general reduction that specifies the properties required for doing so while preserving performance guarantees. In concurrent work, Kolev et al. (2024) combine hybrid training with pessimism for more interaction-efficient model-based IRL, using a model ensemble disagreement auxiliary cost in their practical RL procedure. In contrast, we focus on the benefits hybrid training + expert resets provides for model-based IRL. This allows us to elide intractable pessimism in theory and ensemble-based approximations in practice.
@@ 3.1 A Game-Theoretic Perspective on Inverse RL
We consider a finite-horizon Markov Decision Process (MDP) (Puterman, 2014) parameterized by $\langle\mathcal{S},\mathcal{A},\mathcal{T},r,H\rangle$ where $\mathcal{S}$, $\mathcal{A}$ are the state and action spaces, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ is the transition operator, $r:\mathcal{S}\times\mathcal{A}\rightarrow[-1,1]$ is the reward function, and $H$ is the horizon. In the inverse RL setup, we see trajectories generated by an expert policy $\pi_{E}:\mathcal{S}\rightarrow\Delta(\mathcal{A})$, but do not know the reward function. Our goal is to find a policy that, no matter what reward function we are evaluated under, performs as well as the expert. We cast this problem as a zero-sum game between a policy player and an adversary that tries to pick out differences between expert and learner policies (Syed & Schapire, 2007; Swamy et al., 2021). More formally, we optimize over policies $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})\in\Pi$ and reward functions $f:\mathcal{S}\times\mathcal{A}\rightarrow[-1,1]\in\mathcal{F}_{r}$. We use $\xi=(s_{1},a_{1},r_{1},\ldots,s_{H},a_{H},r_{H})$ to denote the trajectory generated by some policy. For theoretical simplicity, we assume that our strategy spaces ($\Pi$ and $\mathcal{F}_{r}$) are convex and compact, that $\mathcal{F}_{r}$ is closed under negation, and that $r\in\mathcal{F}_{r},\pi_{E}\in\Pi$. Using $J(\pi,\hat{r})=\mathbb{E}_{\xi\sim\pi}[\sum_{h=1}^{H}\hat{r}(s_{h},a_{h})]$ to denote the value of policy $\pi$ under reward function $\hat{r}$, we can express our objective as
@@
Indeed, one may need to explore the entire state space for the optimal policy when given an arbitrary reward function. However, in inverse RL we explicitly choose reward functions that rate the expert higher than the history of learner policies. Thus, the fact that the standard recipe for both primal and dual IRL completely ignore the expert demonstrations during policy optimization seems suboptimal. This begs the question: how can we give the learner examples of expert behavior during policy optimization to reduce the amount of exploration required to find good states? We now discuss the core principle that underlies multiple ways to do so while preserving performance guarantees.
@@ 4 Experiments
Figure 4: We see HyPER and HyPE achieve the highest reward on the MuJoCo locomotion benchmark. Further, the performance gap increases with the difficulty of the environment (i.e. how far right a plot is in the above figure). We run all model-free algorithms for 1 million environment steps. Due to the higher interaction efficiency of model-based approaches, we only run HyPER for 150k environment steps, after which the last reward is extended horizontally across. We compute standard error across 5 seeds for HyPER, and across 10 seeds for all other algorithms.
@@
We implement HyPE by updating the policy and critic networks in Soft Actor Critic (Haarnoja et al., 2018) with expert and learner samples. We implement HyPER by running model-based policy optimization (Janner et al., 2019) and resetting to expert states in the learned model. No reward information is provided in either case, so we also train a discriminator network. We compare against five baselines: BC (behavioral cloning, Pomerleau (1988)), IQLearn (Garg et al., 2021), MM (a baseline moment-matching algorithm that uses an integral probability metric instead of Jensen-Shannon divergence as suggested by Swamy et al. (2022)), BC-Reg (MM with an added Mean-Square Error Loss on the actor update), and FILTER (IRL with resets to expert states,  Swamy et al. (2023)). Appendix B includes additional implementation details and hyperparameters. We release the code we used for all of our experiments at https://github.com/jren03/garage.
MuJoCo Experiments. On the MuJoCo locomotion benchmark environments (Brockman et al., 2016), all learners are provided with 64 demonstration trajectories  [^4] We perform an ablation of all methods considered in the limited demonstration regime in Appendix D.2. from an RL expert to mitigate finite sample concerns. Given simple behavior cloning can match expert performance under these conditions (Swamy et al., 2021), we harden the problem, noise is injected into the environment in the following manner: with probability $p_{tremble}$, a random action is executed instead of the one chosen by the policy. As seen in Figure 4, HyPE and HyPER converge much quicker to expert performance compared to baselines, and the gap increases as environments get more difficult (further right in the figure). While our algorithms have the same worst-case performance bound as FILTER, HyPE and HyPER show to be empirically more sample efficient. We hypothesize this is because HyPE and HyPER use expert states and actions, while FILTER only uses expert states. We find these results to be particularly exciting, as HyPE and HyPER algorithms show competitive performance and sample efficiency without needing expert resets. Finally, IQLearn fails to match BC performance on some environments, and shows signs of unstable training on others. We suspect this is due to the need to model environment stochasticity implicitly.
@@
5 Conclusion
Despite the many benefits interactive approaches to imitation learning like inverse RL provide, they suffer from a high level of interaction complexity due to their reduction to repeated RL. In response, we derive a new paradigm of efficient IRL algorithms via a novel reduction to expert-competitive RL. We then instantiate this reduction with hybrid, deriving both model-based and model-free algorithms, and show that we can dramatically reduce the amount of interaction required to compute a strong policy by utilizing expert demonstrations during policy search. By preserving the benefits of interaction in imitation while reducing unnecessary exploration, we believe that we have provided the proper algorithmic foundations to take advantage of recent developments in hybrid RL and model-based RL on problems of interest in domains like robotics.
Acknowledgments
JR and SC are supported by the NSF RI #2312956. ZSW is supported in part by the NSF FAI Award #1939606, a Google Faculty Research Award, a J.P. Morgan Faculty Award, a Facebook Research Award, an Okawa Foundation Research Grant, and a Mozilla Research Grant. GS would like to thank Yuda Song for many helpful edits and being a hybrid RL oracle throughout the course of this project and Ken Nakamura for comments on proof style.
Impact Statement
@@ Acknowledgments
Contribution Statements
  • JR implemented both algorithms and baselines, performed all experiments and ablations, and released a high-quality open-source implementation.
  • GS proposed the project, derived the key novel reduction, came up with the algorithms, proved all theorems, wrote most of the paper, and provided extensive assistance with debugging all of the experiments.
  • SC came up with the proof of Lemma 3.5 and wrote Section 3.5 of the paper.
  • ZSW, JAB, and SC advised the project.
@@ A.1 Miscellaneous Lemmas
$+\mathbb{E}_{\xi\sim\pi_{E},M^{\star}}\left[\sum_{h}^{H}\mathbb{E}_{\begin{subarray}{c}s^{\star}_{h+1}\sim M^{\star}\\ a\sim\pi(s_{h+1}^{\star})\end{subarray}}[\hat{Q}_{r}(s^{\star}_{h+1},a)]-\mathbb{E}_{\begin{subarray}{c}\hat{s}_{h+1}\sim\hat{M}\\ a\sim\pi(\hat{s}_{h+1})\end{subarray}}[\hat{Q}_{r}(\hat{s}_{h+1},a)]\right]$
$+\mathbb{E}_{\xi\sim\pi,M^{\star}}\left[\sum_{h}^{H}\mathbb{E}_{\begin{subarray}{c}\hat{s}_{h+1}\sim\hat{M}\\ a\sim\pi(\hat{s}_{h+1})\end{subarray}}[\hat{Q}_{r}(\hat{s}_{h+1},a)]-\mathbb{E}_{\begin{subarray}{c}s^{\star}_{h+1}\sim M^{\star}\\ a\sim\pi(s_{h+1}^{\star})\end{subarray}}[\hat{Q}_{r}(s^{\star}_{h+1},a)]\right].$
@@
$=\mathbb{E}_{\xi\sim\pi_{E},M^{\star}}\left[\sum_{h=0}^{H}\left(\mathbb{E}_{\begin{subarray}{c}s^{\star}_{h+1}\sim M^{\star}(s_{h},a_{h})\\ a\sim\pi_{E}(s^{\star}_{h+1})\end{subarray}}[\hat{Q}_{r}(s^{\star}_{h+1},a)]-\mathbb{E}_{\begin{subarray}{c}s^{\star}_{h+1}\sim M^{\star}(s_{h},a_{h})\\ a\sim\pi(s^{\star}_{h+1})\end{subarray}}[\hat{Q}_{r}(s^{\star}_{h+1},a)]\right)\right.$
$\left.-\sum_{h=0}^{H}\left(\mathbb{E}_{\begin{subarray}{c}\hat{s}_{h+1}\sim\hat{M}(s_{h},a_{h})\\ a\sim\pi_{E}(\hat{s}_{h+1})\end{subarray}}[\hat{Q}_{r}(\hat{s}_{h+1},a)]-\mathbb{E}_{\begin{subarray}{c}s^{\star}_{h+1}\sim M^{\star}(s_{h},a_{h})\\ a\sim\pi(s^{\star}_{h+1})\end{subarray}}[\hat{Q}_{r}(s^{\star}_{h+1},a)]\right)\right].$
@@
(T1) + (T3) $=\mathbb{E}_{\xi\sim\pi_{E},M^{\star}}\left[\sum_{h=0}^{H}\left(\mathbb{E}_{\begin{subarray}{c}a\sim\pi_{E}(s_{h})\end{subarray}}[\hat{Q}_{r}(s^{\star}_{h+1},a)]-\mathbb{E}_{\begin{subarray}{c}a\sim\pi(s_{h})\end{subarray}}[\hat{Q}_{r}(s^{\star}_{h+1},a)]\right)\right.$
$\left.+\sum_{h=0}^{H}\left(\mathbb{E}_{\begin{subarray}{c}s^{\star}_{h+1}\sim M^{\star}(s_{h},a_{h})\\ a\sim\pi(s^{\star}_{h+1})\end{subarray}}[\hat{Q}_{r}(s^{\star}_{h+1},a)]-\mathbb{E}_{\begin{subarray}{c}\hat{s}_{h+1}\sim\hat{M}(s_{h},a_{h})\\ a\sim\pi_{E}(\hat{s}_{h+1})\end{subarray}}[\hat{Q}_{r}(\hat{s}_{h+1},a)]\right)\right].$
@@
$+2H^{2}\mathbb{E}_{s,a\sim d_{\tilde{\pi}}}\left[D_{\text{TV}}(M^{\star}(s,a),\hat{M}(s,a))\right],$
@@ B.1 MuJoCo Tasks
Discriminator. For our discriminator, we start with a learning rate of $48e-$ 4 and decay it linearly over outer-loop iterations. For all model-free MuJoCo experiments, we update the discriminator every 10,000 actor steps. For model-based MuJoCo experiments, we update the discriminator every 2,000 actor steps for Hopper, and every 10,000 actor steps for Ant, Humanoid, and Walker.
@@
HyPER. For HyPER, we use the implementation from Pineda et al. (2021) with modifications on the actor update according to Vemula et al. (2023), and turn on the entropy bonus. The model is provided the same demonstration dataset of 64,000 transition tuples as model-free experiments. We use an ensemble of discriminators equal to the model ensemble size, and take the minimum value of the ensemble each forward pass. Further, we find that clipping the discriminator output, adding the same weight decay to the model and actor networks, and using an exponential moving average of the actor weights during inference helps stabilize performance for some environments. We list the specific hyperparameters used for each environment in Tables 3 to 6.
@@ B.2 D4RL Tasks
HyPE. Both HyPE and HyPE+Resets use the same TD3+BC optimizer and hyperparameters for the actor as MM and FILTER, and is pretrained with 10,000 steps of behavioral cloning. HyPE+Resets uses $\alpha=1.0$ to always reset to expert states.
@@ Appendix C Antmaze Model Pretraining
Appendix D Additional Ablations
D.1 Classical Adversarial Imitation Learning Methods
Our inverse RL baseline MM can be thought of as a significantly improved variant of the classical methods such as GAIL (Ho & Ermon, 2016). More explicitly, we implement the four changes to GAIL suggested by (Swamy et al., 2022) which, when combined, lead to a significant improvement over off-the-shelf implementations of GAIL. Explicitly, these are 1) using a Wasserstein GAN loss rather than the original GAN loss, 2) using SAC instead of PPO, 3) using gradient penalties in the loss function, and 4) using the Optimistic Adam Optimizer. We would like to point readers to Appendix C of Swamy et al. (2022) for ablations on each of these components. To ablate the joint benefits of these modifications, we compare our MM baseline to GAIL (as implemented in https://github.com/ikostrikov/pytorch-a2c-ppo-acktr-gail, a popular implementation with more than 3.5k Github stars) on the Humanoid environment, and report the average performance over 10 seeds. Figure 7 shows that while GAIL achieves some improvement over 1 million environment interactions, it massively under performs MM, reaffirming the fact that MM is a very strong baseline for standard IRL in and of itself.
Figure 7: Comparison of our MM baseline against GAIL (Ho & Ermon, 2016), reporting the average and standard error over 10 seeds. While GAIL make gradual improvement over 1 million environment interactions, it is far from the performance of MM.
D.2 Low Data Regime
We present in Figure 8 an ablation of the performance of various algorithms in the low-data regime. Specifically, we take just 5 trajectories to train each algorithm, as opposed to the 64 used in the main experiments. We observe that HyPE still has the quickest convergence over other model-free IRL baselines, even in the low-data regime. HyPER performs rather poorly in the low-data regime, which we hypothesize is due to the difficulty of learning an accurate model from such limited data. However, given a world model can be learned from other data sources which might be more plentiful in practice (e.g. suboptimal demonstrations, multi-task demonstrations, and even robot play data), we believe there are multiple remedies to this issue in practice.
Figure 8: Mean and standard error of various algorithms when provided only 5 expert demonstrations over 10 seeds. HyPE still outperforms existing baselines. HyPER suffers a performance drop, which is expected as low amounts of expert data may not be sufficient for fitting a good model. We believe there are various ways to improve model fitting in practice in the presence of few expert demonstrations, such as multi-task or robot play data.